# Spatial Disparities in Fire Shelter Accessibility: Capacity Challenges in the Palisades and Eaton Fires

### v1

@@ Abstract
The increasing frequency and severity of wildfire in California, exacerbated by prolonged drought and environmental changes, pose significant challenges to urban community resilience and equitable emergency response. The study investigates issues of accessibility to shelters during the Palisades and Eaton Fires which started in January 2025 in Southern California that led to over 180,000 displacements and the loss of 16,000 structures. Despite coordinated efforts of many organizations’ emergency assistance, shelter shortages left many evacuees without safety or accessible refuge. This research aims to measure shelter accessibility during the fires’ peak, evaluate whether existing shelter capacity met the demand, and identify spatial disparities in access. Results reveal severe shelter shortages and pronounced inequities in access to shelters, particularly in geographically isolated regions and mountainous areas. Our simulations of shelter placement strategies using a capacity-based algorithm and a proximity-based approach demonstrate potential improvements in both shelter accessibility and equitable access to shelters. The findings underscore the critical need for strategic shelter planning and infrastructure development to enhance disaster readiness and reduce vulnerability in regions that frequently experience wildfires.
@@
spatial accessibility , wildfire , disparities , shelter , simulation
@@
[…]
[usc]organization=Spatial Sciences Institute, University of Southern California, addressline=Crocker Plaza, 3616 Trousdale Parkway, city=Los Angeles, postcode=90089, state=CA, country=USA
@@ 1 Introduction
Although significant challenges to shelter accessibility arose during the Palisade and Eaton Fires, spatial disparities in shelter availability have not been thoroughly examined. Hence, this research aims to investigate shelter accessibility during the crisis while assessing whether they provided sufficient capacity to accommodate evacuees. Additionally, it examines whether evacuees had equitable access to shelters in different regions, while also exploring how geographic differences in shelter accessibility were shaped by traffic congestion during the fire events. Furthermore, this research suggests potential new shelter locations that could improve overall accessibility and ensure more equitable shelter access throughout all evacuation zones. Results and findings from this research can assist shelter placement and distribution of critical resources to better support for displaced populations during future wildfire emergencies.
This research is intended to answer research questions below.
@@
    How effective was shelter accessibility during the peak of the Palisades and Eaton Fires, and did the total shelter capacity adequately support the displaced population?
@@
    What were the travel times to the nearest shelters for evacuees, and how did traffic congestion impact spatial accessibility and affected disparities in shelter access?
@@
    How did shelter accessibility vary across different regions, and to what extent did spatial disparities exist in terms of equitable access to shelters?
@@
    How can the strategic placement of new shelters improve overall accessibility and ensure equitable shelter access across all regions during disasters?
@@ 2 Related Work
Research shows that social and geographic inequalities in access to emergency shelters limit disaster response effectiveness and place vulnerable populations at greater risk. Ermagun et al. (2024) noted that the elderly, individuals with disabilities, and Hispanic communities are at greater risk due to limited shelter access [26]. Moreover, geographic disparities create extra challenges for suburban and rural areas, which often struggle with sparse and insufficient emergency facilities [26, 27, 28].
@@
Although studies on emergency shelters’ spatial accessibility have increased, there are still few that are especially focused on wildfires. For example, Ermagun and Janatababi [26] measured accessibility to wildfire shelters but did not account for the dynamic interactions among supply (shelters), demand (population), and travel time–key factors incorporated in the Floating Catchment Area (FCA) framework. Furthermore, the majority of current studies evaluate spatial accessibility to urban emergency facilities such as shelters, fire stations, and emergency centers (e.g., [39, 45, 46, 47, 48]), while often overlooking spatial accessibility in wildfire-prone regions. The populations in these areas are typically scattered throughout forested landscapes and have limited road infrastructure, which complicates emergency response efforts and shelter accessibility. Filling these research gaps will play a crucial role in enhancing wildfire disaster preparedness and ensuring equitable shelter access.
@@ 3.1 Shelter Capacity
We gathered the locations, addresses, and overnight counts of available shelters as of January 12, 2025, from the Cal OES News website [50]. The capacity of shelters open on January 12, 2025, as well as other potential shelters used in this study, was collected from the website, National Shelter System Facilities (NSSF) [51]. The dataset includes 70,317 facilities across the U.S. that can accommodate individuals during an evacuation. While the downloaded shelter data included evacuation capacity information, this information was not consistently available for all shelters. Even though there were 11 shelters that remained open throughout the active Palisades and Eaton fires (as shown in Table 1), capacity information was available for only six of them. For the remaining five shelters, we estimated capacity by assessing the available indoor space using various online sources. For instance, the Van Nuys/Sherman Oaks Recreation Center was found to have approximately 15,000 square feet of space, based on information obtained from a website [52]. We then calculated the usable shelter space according to FEMA guidelines.
FEMA’s Guidelines for Mass Care Shelter Operations (FEMA P-785) specify that non-occupiable areas must be accounted for when planning emergency shelter space [53]. Similarly, the Sphere Handbook implies that only about 65–75% of the total indoor area is usable for sheltering, with approximately 25–35% allocated to non-occupiable spaces such as hallways, mechanical rooms, and storage areas [54]. Based on these recommendations, we assumed that 70% of each facility’s total space was suitable for sheltering. Additionally, we assumed that each person required 100 square feet to ensure a desirable shelter environment. This estimate is based on CDC guidelines, which recommend a minimum of 60 square feet per person in evacuation shelters [55]. However, FEMA further advises that shelters should provide one toilet for every 20 people and one shower for every 25 people. Since no detailed information was available regarding the number of toilets or showers in each shelter facility, we adopted a more conservative estimate of 100 square feet per person to account for potential limitations. A detailed breakdown of how capacity was estimated for each of the five shelters is provided in A. Table 1 shows all eleven shelter locations and capacities during the Palisades and Eaton wildfires. For facilities where capacity information is available on the website [51], the value is recorded in the capacity column. For facilities where capacity data is unavailable, the estimated capacity is provided in a separate column as estimated capacity.
@@
Van Nuys/Sherman Oaks Recreation Center 100
@@ 3.3 Road Network
The road data were collected from OpenStreetMap (OSM) [60], which are a widely used dataset to compute travel cost in previous accessibility studies [61, 58, 35, 34, 33]. We retrieved the OSM road network dataset using the Python package OSMnx [62], which allows for the downloading and analysis of street networks from OpenStreetMap. Given the sparse road networks, particularly in mountainous regions, we included all drivable roads, including service roads such as alleys and parking lot roads, to ensure comprehensive coverage. Figure 3 depicts the road network used in our study, which is extensive enough to cover wildfire-related evacuation routes surrounding shelters. The final road network dataset we downloaded contained 701,470 line features.
Figure 3: Road network data retrieved from OpenStreetMap (OSM) for our study area.
@@ 4.1 Accessibility Measurement
We applied the Enhanced Two-Step Floating Catchment Area (E2SFCA) method to assess spatial accessibility between supply locations (i.e., shelters) and demand (i.e., population within each grid cell located in evacuation and warning zones). The E2SFCA method [32] builds upon the original Two-Step Floating Catchment Area (2SFCA) method developed by [37]. It has been widely adopted in numerous studies, particularly for evaluating accessibility to healthcare services [35, 34, 33]. The E2SFCA method calculates accessibility in two steps, considering supply (shelter capacity), demand (population in each grid cell), and the travel cost between supply and demand locations, measured in time or distance. A key feature of this method is its incorporation of distance decay, which assumes that locations closer to a supply point have a greater influence, whereas those farther away have a lesser impact[63]. As the name suggests, the method consists of two steps.
@@
    $t_{0}$ is the travel time threshold, set to 2 hours in this study.
@@
The Spatial Access module in PySAL (The Python Spatial Analysis Library) [64] includes an implementation of the Enhanced Two-Step Floating Catchment Area (E2SFCA) method, available through the function Access.enhanced_two_stage_fca [65]. One of the input parameters required for this function is the weight function $W_{ij}$, as defined in the above equations. Users can specify either a step function or a Gaussian weight function with a specified width ($\sigma$). In this study, we applied a Gaussian weight function with $\sigma$ = 30.
@@ 4.2 Computing Travel Cost
Travel time was computed based on the length and maximum speed of road segments within street networks extracted from OpenStreetMap (OSM). As indicated in the equation above, the accessibility score is influenced not only by supply and demand but also by travel cost, which plays a crucial role in shaping access. The OSM street network includes road classification, segment length, and maximum speed, all of which vary based on road type. For instance, motorways, secondary roads, and service roads have maximum speeds of 65 mph, 35 mph, and 15 mph, respectively. When the maximum speed was unavailable for a road segment, we imputed a value based on the mean maximum speed value for edges categorized by highway type, using a function provided by the OSMnx Python package [62]. Using the same package, we also computed travel times by leveraging the maximum speed and segment length for each road or line segment. Additionally, to compute the shortest paths from each demand location to all shelter locations, we employed Dijkstra’s algorithm, a single-source shortest path algorithm, where the weight of an edge in the network is an estimated travel time from the starting vertex to the ending vertex of the edge. The travel cost, essential for measuring accessibility, was estimated by computing travel times along the shortest path from each demand location (i.e., each grid cell in Figure 2) to each shelter location. In addition, when computing spatial accessibility under traffic congestion [66], a maximum travel speed of 10 kph was assigned to the road network within a five-kilometer buffer zone around the Palisades and Eaton evacuation areas shown in Figure 2. This assumes that vehicles could not travel at the designated road speeds and are instead limited to an average of 10 kph due to severe traffic jams and gridlock [66].
@@ 4.4 Workflow of This Study
The study’s workflow is organized into four distinct cases, each tailored to evaluate accessibility, as detailed in Table 2, under varying spatial and situational conditions. Specifically, we considered different configurations of shelter (i.e., supply) availability, including hypothetical shelters; diverse demand areas based on evacuation orders, warning zones, and the presence or absence of active fire areas; and variations in travel cost, accounting for traffic congestion and road closures.
Cases 1 through 3 focus on real-world conditions during the peak wildfire crisis on January 12, while Case 4 examines hypothetical scenarios aimed at enhancing shelter access. On this day, a significant gap existed between available shelter capacity and the number of people living in evacuation zones. The eight operational shelters available on January 12 (the top eight listed in Table 1) had a combined capacity of only 5,224 people. In contrast, population estimates from LandScan Global 2023 indicate that 44,348 individuals were located in evacuation order zones and 42,263 in warning zones, totaling 86,611 people at risk. This left approximately 81,387 people without access to designated shelters. The large gap between available shelter capacity and evacuation demands highlights the immediate need to increase shelter infrastructure to better protect evacuees during such emergencies.The critical gap between available shelter spaces and evacuation requirements on January 12 led to the creation of Case 4, which investigated potential solutions by adding hypothetical shelter sites. Specifically, we examined potential shelter sites drawn from the National Shelter System Facilities dataset ([51]), which includes buildings that could potentially serve as shelters and specify their capacities.
The objective of Case 4 is to identify the locations and capacities of additional shelters needed to accommodate the entire evacuated population and to improve overall accessibility. To achieve this, we designed two sub-scenarios that determine which shelters should be selected and how they should be spatially distributed.
  • 1.
@@
    Case 4.1: Capacity-based Approach — The idea of this approach is that when there are too many shelters with small capacities, resources needed for evacuees–such as cots, blankets, food, and medical assistance–must be distributed across multiple locations. By prioritizing larger shelters, the approach aims to minimize logistical efforts and enhance the efficiency of resource allocation during emergencies. Our capacity-based greedy algorithm consists of two steps: filtering and refinement. In the filtering step, starting from the demand areas, the algorithm incrementally searches for the nearest shelters and adds them one by one until their combined capacity exceeds k times the required demand. In our empirical analysis, when k=2, shelters within a reasonable distance are included. In the refinement step, the algorithm selects shelters in descending order of capacity from the candidates, starting with the largest, until the total capacity becomes sufficient to accommodate the estimated population in need.
  • 2.

### v2

@@ Abstract
The increasing frequency and severity of wildfire in California, exacerbated by prolonged drought and environmental changes, pose significant challenges to urban community resilience and equitable emergency response. The study investigates issues of accessibility to shelters during the Palisades and Eaton Fires which started in January 2025 in Southern California that led to over 180,000 displacements and the loss of 16,000 structures. Despite coordinated efforts of many organizations’ emergency assistance, shelter shortages left many evacuees without safety or accessible refuge. This research aims to measure shelter accessibility during the fires’ peak, evaluate whether existing shelter capacity met the demand, and identify spatial disparities in access. Findings reveal severe shelter shortages and pronounced inequities in access to shelters, particularly in geographically isolated regions and mountainous areas. To address these challenges, we implemented shelter placement strategies using both capacity-based and distance-based approaches, demonstrating potential improvements in accessibility and equity. The findings underscore the critical need for strategic shelter planning and infrastructure development to enhance disaster readiness and reduce vulnerability in regions that frequently experience wildfires.
@@
spatial accessibility , wildfire , disparities , shelter
@@ 1 Introduction
This study focuses specifically on the recent Palisades and Eaton wildfires, which exposed significant challenges in shelter accessibility during large-scale evacuations. Despite the severity of these events, spatial disparities in access to wildfire shelters have not been systematically examined. Accordingly, this research aims to evaluate the spatial accessibility of shelters during the Palisades and Eaton Fires, assessing whether the available facilities provided sufficient capacity to accommodate evacuees. It also investigates the equity of shelter access across different regions and analyzes how geographic variations in accessibility were influenced by traffic congestion during the evacuations. Furthermore, the study identifies potential new shelter locations to enhance overall accessibility and promote more equitable shelter distribution across all evacuation zones. Findings from this research will inform future shelter siting and emergency planning, supporting more effective protection and resource allocation for displaced populations during future wildfire disasters.
This research is intended to answer the research questions below.
@@
    To what extent was shelter accessibility effective during the peak of the Palisades and Eaton Fires—two of the most destructive wildfires that occurred in close proximity to densely populated residential areas—and did the available shelter capacity adequately accommodate the large number of displaced residents?
@@
    What were the travel times to the nearest shelters for evacuees during the Palisades and Eaton Fires, and how did severe traffic congestion and road network disruptions influence spatial accessibility and exacerbate disparities in reaching safe shelters?
@@
    How did shelter accessibility vary across the affected regions of the Palisades and Eaton Fires, and to what extent did spatial disparities emerge in terms of equitable access to emergency shelters among nearby residential communities?
@@
    How could the strategic placement of additional shelters within the fire-affected areas improve overall accessibility and promote equitable, timely shelter access for residents living in high-risk and densely populated zones during wildfire emergencies such as the Palisades and Eaton Fires?
@@ 2 Related Work
2.2 Social and Geographic Inequities in Shelter Accessibility
Research shows that social and geographic inequalities in access to emergency shelters limit disaster response effectiveness and place vulnerable populations at greater risk. Ermagun et al. noted that the elderly, individuals with disabilities, and Hispanic communities are at greater risk due to limited shelter access [23]. Moreover, geographic disparities create extra challenges for suburban and rural areas, which often struggle with sparse and insufficient emergency facilities [23, 84, 51].
@@ 2.2 Social and Geographic Inequities in Shelter Accessibility
2.3 Spatial Accessibility and the Floating Catchment Area (FCA) Framework
@@
2.4 Research Gaps in Wildfire Shelter Accessibility
Although studies on emergency shelters’ spatial accessibility have increased, there are still few that are especially focused on wildfires. For example, Ermagun and Janatababi [23] measured accessibility to wildfire shelters but did not account for the dynamic interactions among supply (shelters), demand (population), and travel time–key factors incorporated in the FCA framework. Furthermore, the majority of current studies evaluate spatial accessibility to urban emergency facilities such as shelters, fire stations, and emergency centers (e.g., [89, 88, 77, 50, 42]), while often overlooking spatial accessibility in wildfire-prone regions. The populations in these areas are typically scattered throughout forested landscapes and have limited road infrastructure, which complicates emergency response efforts and shelter accessibility. Filling these research gaps will play a crucial role in enhancing wildfire disaster preparedness and ensuring equitable shelter access.
@@ 3.1 Shelter Capacity
We gathered the locations, addresses, and overnight counts of available shelters as of January 12, 2025, from the Cal OES News website [13]. The capacity of shelters open on January 12, 2025, as well as other potential shelters used in this study, was collected from the website, National Shelter System Facilities (NSSF) [35]. The dataset includes 70,317 facilities across the U.S. that can accommodate individuals during an evacuation. While the downloaded shelter data included evacuation capacity information, this information was not consistently available for all shelters. Even though there were 11 shelters that remained open throughout the active Palisades and Eaton fires (as shown in Table 1), capacity information was available for only six of them. For the remaining five shelters, we estimated capacity by assessing the available indoor space using various online sources. For instance, the Van Nuys/Sherman Oaks Recreation Center was found to have approximately 15,000 square feet of space, based on information obtained from a website [4]. We then calculated the usable shelter space according to the Federal Emergency Management Agency (FEMA) guidelines.
According to FEMA’s Guidelines for Mass Care Shelter Operations (FEMA P-785), planners must exclude non-occupiable areas—such as restrooms, kitchens, hallways, and administrative spaces—when calculating available shelter capacity [1]. Following these recommendations, we assumed that approximately 70% of each facility’s total floor area was suitable for occupancy. Furthermore, to ensure a comfortable and safe shelter environment, we allocated 100 square feet per person. This estimate exceeds the Centers for Disease Control and Prevention or CDC’s shelter assessment guidance, which recommends a minimum of 60 square feet per person in evacuation shelters [16]. However, FEMA further advises that shelters should provide one toilet for every 20 people and one shower for every 25 people. Since no detailed information was available regarding the number of toilets or showers in each shelter facility, we adopted a more conservative estimate of 100 square feet per person to account for potential limitations. A detailed breakdown of how capacity was estimated for each of the five shelters is provided in Section A. Table 1 shows all eleven shelter locations and capacities during the Palisades and Eaton wildfires. For facilities where capacity information is available on the website [35], the value is recorded in the capacity column. For facilities where capacity data is unavailable, the estimated capacity is provided in a separate column as estimated capacity.
@@
Van Nuys/Sherman Oaks Recreation Center 105
@@ 3.3 Road Network
The road data were collected from OpenStreetMap (OSM) [31], which are a widely used dataset to compute travel cost in previous accessibility studies [44, 49, 39, 41, 60]. We retrieved the OSM road network dataset using the Python package OSMnx [9], which allows for the downloading and analysis of street networks from OpenStreetMap. Given the sparse road networks, particularly in mountainous regions, we included all drivable roads, including service roads such as alleys and parking lot roads, to ensure comprehensive coverage. The road network downloaded in our study, as shown in Section B, is extensive enough to cover wildfire-related evacuation routes surrounding shelters. The final road network dataset we downloaded contained 701,470 line features.
@@ 4.1 Accessibility Measurement
We applied the E2SFCA method to assess spatial accessibility between supply locations (i.e., shelters) and demand (i.e., population within each grid cell located in evacuation and warning zones). The E2SFCA method [47] builds upon the original 2SFCA method developed by [48]. It has been widely adopted in numerous studies, particularly for evaluating accessibility to healthcare services [39, 41, 60]. The E2SFCA method calculates accessibility in two steps, considering supply (shelter capacity), demand (population in each grid cell), and the travel cost between supply and demand locations, measured in time or distance. A key feature of this method is its incorporation of distance decay, which assumes that locations closer to a supply point have a greater influence, whereas those farther away have a lesser impact[59]. As the name suggests, the method consists of two steps.
@@
    $t_{0}$ is the travel time threshold, set to 2 hours in this study. See the Section C for the rationale behind selecting the 2-hour threshold.
@@
  • 5.
    $t_{0}$ is the travel time threshold
The Spatial Access module in PySAL (The Python Spatial Analysis Library) [62] includes an implementation of the E2SFCA method, available through the function Access.enhanced_two_stage_fca [61]. One of the input parameters required for this function is the weight function $W_{ij}$, as defined in the above equations. Users can specify either a step function or a Gaussian weight function with a specified width ($\sigma$). In this study, we applied a Gaussian weight function with $\sigma$ = 30. The rationale for selecting this value is discussed in Section C.
@@ 4.2 Computing Travel Cost
Travel time was computed based on the length and maximum speed of road segments within street networks extracted from OSM. As indicated in the equation above, the accessibility score is influenced not only by supply and demand but also by travel cost, which plays a crucial role in shaping access. The OSM street network includes road classification, segment length, and maximum speed, all of which vary based on road type. For instance, motorways, secondary roads, and service roads have maximum speeds of 65 mph, 35 mph, and 15 mph, respectively. When the maximum speed was unavailable for a road segment, we imputed a value based on the mean maximum speed value for edges categorized by highway type, using a function provided by the OSMnx Python package [9]. Using the same package, we also computed travel times by leveraging the maximum speed and segment length for each road or line segment. Additionally, to compute the shortest paths from each demand location to all shelter locations, we employed Dijkstra’s algorithm, a single-source shortest path algorithm, where the weight of an edge in the network is an estimated travel time from the starting vertex to the ending vertex of the edge. The travel cost, essential for measuring accessibility, was estimated by computing travel times along the shortest path from each demand location (i.e., each grid cell in Figure 2) to each shelter location. In addition, when computing spatial accessibility under traffic congestion [46], a maximum travel speed of 10 kph was assigned to the road network within a five-kilometer buffer zone around the Palisades and Eaton evacuation areas shown in Figure 2. This assumes that vehicles could not travel at the designated road speeds and are instead limited to an average of 10 kph due to severe traffic jams and gridlock [46].
@@ 4.4 Workflow of This Study
The study workflow is structured into four distinct cases, each designed to evaluate accessibility under different spatial and situational conditions (Table 2). Across these cases, we vary three key components: shelter availability (supply)—including both operational and hypothetical shelters; demand definitions, based on evacuation orders, warning zones, and the inclusion or exclusion of areas affected by active fires; and travel costs, accounting for wildfire-induced traffic congestion and road closures.
Cases 1 through 3 represent real-world conditions during the peak of the wildfire crisis on January 12. As an illustrative example, Case 3 in Figure 3 corresponds to process P7, which models an event-driven scenario that departs from baseline accessibility assumptions by explicitly incorporating wildfire-related disruptions. Process P7 has two primary inputs: supply (I3), representing operational shelters, and demand (I8), defined as the population located within evacuation warning zones while excluding areas impacted by active fires. This exclusion is based on the assumption that residents do not remain in areas that are already burned or actively burning. Consequently, the demand region is constructed by erasing the fire perimeter zones (I5) from the evacuation warning zones (I6). The I6 layer is generated in P1 by clipping LandScan Global 2023 population estimates at a 1 km resolution (I1) with evacuation order and warning zones associated with the Palisades and Eaton wildfires (I2).
Process P7 further incorporates I7, which represents surrounding regions presumed to experience increased traffic congestion during the wildfire event. These regions are created in P2 by applying a 5 km buffer around evacuation warning zones derived from I2. Within these buffered areas, travel speed is assumed to be reduced to 10 kph, reflecting wildfire-related congestion. Together, P7 computes travel times by assuming no road access within fire perimeters and reduced travel speeds in surrounding wildfire-affected zones, enabling the estimation of shelter accessibility under conditions of road closures and severe traffic congestion. The rationale for selecting a 5 km buffer to delineate congestion zones and a travel speed of 10 kph is discussed in detail in Section D.
In contrast, Case 4 examines hypothetical scenarios aimed at improving shelter accessibility under extreme evacuation demand. On January 12, a substantial mismatch existed between available shelter capacity and the population residing in evacuation zones. The eight operational shelters available on that day (the top eight listed in Table 1) had a combined capacity of 5,224 people, whereas LandScan Global 2023 estimates indicate that 44,348 individuals were located in evacuation order zones and 42,263 in warning zones, totaling 86,611 people at risk. As a result, approximately 81,387 individuals lacked access to designated shelters. This pronounced disparity between shelter capacity and evacuation demand underscores the urgent need to expand shelter infrastructure during large-scale wildfire emergencies.
Motivated by this critical gap, Case 4 explores potential mitigation strategies by introducing hypothetical shelter sites. Candidate locations are drawn from the National Shelter System Facilities dataset [35], which identifies buildings that could potentially serve as shelters and provides information on their capacities.
Case 4.1: Capacity-based Approach
@@
— The idea of this approach is that when there are too many shelters with small capacities, resources needed for evacuees (such as cots, blankets, food, and medical assistance) must be distributed across multiple locations. By prioritizing larger shelters, the approach aims to minimize logistical efforts and enhance the efficiency of resource allocation during emergencies. Our capacity-based greedy algorithm consists of two steps: filtering and refinement. In the filtering step, starting from the demand areas, the algorithm incrementally searches for the nearest shelters and adds them one by one until their combined capacity exceeds $k$ times the required demand. In our empirical analysis, when $k=2$, shelters within a reasonable distance are included. In the refinement step, the algorithm selects shelters in descending order of capacity from the candidates, starting with the largest, until the total capacity becomes sufficient to accommodate the estimated population in need.
Input: Set of demand areas $D$
Input: Set of shelters $S$ with capacities $C(s)$ for each $s\in S$
Input: Population demand $P_{d}$ for each $d\in D$
Input: Multiplier $k$
Output: Set of selected shelters $S_{\text{selected}}$
1
$S_{\text{selected}}\leftarrow\emptyset$ ;
// Initialize the final list of selected shelters
2
3foreach demand area $d\in D$ do
 $S_{\text{candidate}}\leftarrow\emptyset$ ;
 // Temporary list of nearby shelters
 $C_{\text{total}}\leftarrow 0$ ;
 // Track cumulative capacity
4
 // Filtering step: prioritize nearby shelters
5   Sort shelters $S$ by distance to $d$ in ascending order;
6
7 foreach shelter $s\in S$ (in sorted order) do
8      Add $s$ to $S_{\text{candidate}}$;
9    $C_{\text{total}}\leftarrow C_{\text{total}}+C(s)$;
10    if $C_{\text{total}}\geq k\cdot P_{d}$ then
       break ;
       // Stop once buffer capacity (k × demand) is met
11
12
13
 // Refinement step: prioritize larger shelters
14   Sort $S_{\text{candidate}}$ by capacity $C(s)$ in descending order;
 $C_{\text{refined}}\leftarrow 0$ ;
 // Track capacity in refined selection
15
16 foreach shelter $s\in S_{\text{candidate}}$ do
17      Add $s$ to $S_{\text{selected}}$;
18    $C_{\text{refined}}\leftarrow C_{\text{refined}}+C(s)$;
19    if $C_{\text{refined}}\geq P_{d}$ then
       break ;
       // Stop once actual demand is met
20
21
22
23
return $S_{\text{selected}}$ ;
// Return selected shelters across demand areas
24
Algorithm 1 Capacity-Based Shelter Selection
Algorithm 1 shows the detail of the capacity-based approach as pseudo code. The algorithm begins by initializing an empty set to store the final selected shelters (Line 1). For each demand area (Line 2), a candidate list of nearby shelters is initialized (Line 3), and a total capacity counter is set to zero (Line 4). In the filtering step (Lines 5–10), the algorithm sorts all shelters by their distance to the current demand area (Line 5), and iteratively adds them to the candidate list until the combined capacity exceeds $k\cdot P_{d}$ (Lines 6–10). This buffer-based threshold ensures that a sufficient number of shelters are considered without exceeding a reasonable distance. Once the candidate set is built, the refinement step begins by sorting it in descending order of shelter capacity (Line 11). Then, starting from the largest shelters, the algorithm selects shelters into the final output set until the actual population demand is met (Lines 12–16). Finally, the set of all selected shelters across all demand areas is returned (Line 17).
Case 4.2: Distance-based Approach